nine books are to be bought by a student.Some cost $6 each and the remainder cost $6.50 each. if the total amount spent was $56, how many books at each price are bought?
Answer by Cromlix(4381)

nine books are to be bought by a student
x + y = 9.....1
Some cost $6 each and the remainder cost $6.50 each. if the total amount spent was $56,
6x + 6.50y = 56.....2
....................
x + y = 9.....1
6x + 6.50y = 56 ......2
Multiply (1) by 6.50
6.50x + 6.50y = 58.50 ...3
6x + 6.50y = 56.....2
Subtract (2) from (3)
0.50x = 2.50
x = 5
Substitute x = 5 into (1)
x + y = 9
5 + y = 9
y = 9 - 5
y = 4.
5 x $6 books + 4 x $6.50 books
Hope this helps.
:-)